OpenStudy (abbysmith12):

@jordanloveangel @_Bugzbunny_ @Preetha @coolaidpower @bohotness @demonchild99 @Nnesha @TheRaggedyDoctor

OpenStudy (anonymous):

need help?

OpenStudy (anonymous):

Please post your question first

OpenStudy (anonymous):

And subsequently tag someone , that way you can get the most help

OpenStudy (abbysmith12):

i tried my laptop was loading

Nnesha (nnesha):

combine like terms

OpenStudy (anonymous):

Basically you have to combine like terms. Do you know how to accomplish this?

OpenStudy (abbysmith12):

no

Nnesha (nnesha):

add number with same variable and same exponent together and number without variable add them tghter

Nnesha (nnesha):

no???

OpenStudy (abbysmith12):

what is that a bad thing

OpenStudy (abbysmith12):

@ScarlettGrace<3

Nnesha (nnesha):

lets do first one together here 3x and x both have x variable and same exponent so you can combine them what is 3x+x = ??

OpenStudy (abbysmith12):

umm 3?

OpenStudy (abbysmith12):

@jordanloveangel

OpenStudy (anonymous):

Like terms are numbers that have the same ending (variable or no variable and same exponent) ex: \[\Large \text{1, 5, and 8 are like terms}\] \[\Large \text{a, 2a, 6a, and 11a are like terms}\] \[\Large x^2,~ 3x^2,~ 5x^2 \text{are like terms}\]

OpenStudy (abbysmith12):

@Nnesha

OpenStudy (anonymous):

You can add or subtract them together like basic numbers, but you have to make sure you keep their endings so \[\Large 3a+2a=5a\]

Nnesha (nnesha):

x have coefficient which is one so its 3x + 1x = ???

OpenStudy (abbysmith12):

4x

Nnesha (nnesha):

yes right now combine number whic doesn't have any variable